Find the value of x in degree 5 (sin ^2)x + 12 sin x + 4 = 0
Answer by stanbon(75887)

Find the value of x in degree 5 (sin ^2)x + 12 sin x + 4 = 0
----------
5sin^2(x) + 12sin(x) + 4 = 0
--------
Factor:
5sin^2(x) + 10sin(x)+2sin(x) + 4 = 0
-----
5sin(x)[sin(x)+2][2((sin(x)+2)] = 0
-----
(sin(x)+2)(5sin(x)+2] = 0
-------------------------------
sin(x) = -2 (extraneous)
sin(x) = -2/5 = -0.4
=========================
Ans: x = arcsin(-0.4) = -23.58 degrees or 180+23.58 degrees
===============================================
Cheers,
Stan H.